OpenStudy (latin749):

⦁ The manager of a fish store has water that is 10% salt and water that is 25% salt. He needs to fill an aquarium with 5 gallons of water that is 20% salt. ⦁ Write a system of linear equations that you can use to determine how many gallons of each type of salt water the manager should combine. Be sure to define your variables. ⦁ Solve the system and determine how many gallons of each type of salt water the manager should combine. Show all your work.

OpenStudy (latin749):

@TheSmartOne

OpenStudy (latin749):

@mathstudent55 @phi @zepdrix

OpenStudy (latin749):

@BGrg007

OpenStudy (phi):

There is a pattern for this type of problem the first thing is to define two variables for the amount of each type. For example, x = # of gallons of 10% salt water y = # of gallons of 25% salt water

OpenStudy (latin749):

Yes that is what I first thought but I didn't know how to setup the equation

OpenStudy (phi):

if you use x gallons of the one type and y gallons of the other type how many gallons total are you using ? any idea ?

OpenStudy (latin749):

x + y ?

OpenStudy (phi):

yes now notice this He needs to fill an aquarium with 5 gallons of water how much do you want x+y to equal ?

OpenStudy (latin749):

5

OpenStudy (phi):

yes. can you write an equation showing that ?

OpenStudy (latin749):

so x + y = 5

OpenStudy (phi):

yes. now we need one more equation. we do that by figuring how much salt we have look at 5 gallons of water that is 20% salt. any idea how much salt that is ?

OpenStudy (latin749):

1 gallon

OpenStudy (phi):

yes, or 5 gallons * 0.20 (we need to know that "formula" to write how much salt is in the x gallons of 10% salt) any idea ?

OpenStudy (latin749):

5 times 0.10?

OpenStudy (latin749):

oh

OpenStudy (phi):

almost, but x is not 5 gallons. x is some # of gallons (which we don't know)

OpenStudy (latin749):

x times 0.10

OpenStudy (phi):

yes and use the same idea for how much salt is in the y gallons (of 25% salt)

OpenStudy (latin749):

so \[x \times 0.10\] and \[y \times 0.25\]

OpenStudy (phi):

and the sum of that salt must equal 5*0.2 = 1

OpenStudy (latin749):

ohhhhhh

OpenStudy (latin749):

\[0.10x + 0.25y = 1 \]

OpenStudy (phi):

yes, now you have two equations and two unknowns.

OpenStudy (latin749):

\[x + y = 5\]

OpenStudy (latin749):

Ok, now I have to solve

OpenStudy (phi):

if you mulitply the 0.1x+0.25y= 1 by 100 (on both sides, and all terms) we get 10x+25y= 100 which gets rid of the decimals now you have 10x+25y= 100 x+ y= 5 if you multiply the bottom equation by -10 (on both sides, all terms) what do you get ?

OpenStudy (latin749):

-10x -10y = -50

OpenStudy (latin749):

So eliminate?

OpenStudy (phi):

yes, now add the two equations the 10x -10x will cancel

OpenStudy (latin749):

15y = 50

OpenStudy (phi):

yes, now divide both sides by 15 on the left 15/15 is 1

OpenStudy (latin749):

y = 3.3?

OpenStudy (latin749):

wait

OpenStudy (latin749):

yeah y = 3.3

OpenStudy (latin749):

x + 3.3 = 5 x = 1.7

OpenStudy (phi):

I would use 10/3 or 3 and 1/3 rather than a decimal

OpenStudy (latin749):

Because of the perpetual 3?

OpenStudy (phi):

yes

OpenStudy (latin749):

y = 3 and 1/3 x = 1 and 2/3

OpenStudy (phi):

yes, that is the answer

OpenStudy (latin749):

Thank you @phi I really appreciate your help, I couldn't get my head around this one

OpenStudy (phi):

yw

OpenStudy (latin749):

Will you be on more today?

OpenStudy (latin749):

I have to brb for 5min but I might need more help as I'm going through these hard lessons

OpenStudy (phi):

ok